OpenStudy (anonymous):

Which equation represents the problem? The circumference of a circle equals the product of two, pi, and the radius. Use C for circumference, 3.14 for pi, and r for the radius

OpenStudy (anonymous):

C = 2 * 3.14 * r or C = 6.28 * r